# Semiformal method I. Pólya’s theorem for the complete graph $K_{\mathbb{N}}$ with complex edge weights

### v6

@@ 2. Indeed, the sum
$\sum_{n=0}^{\infty}h(W(n))x^{n}=\frac{1}{1-\xi x}\,.$}
In particular, for $\xi=0$ we have $\sum_{n=0}^{\infty}h(W(n))x^{n}=1$.
Proof. Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be light and $\xi$-convex. We prove by induction on $n\in\mathbb{N}_{0}$ that $h(W(n))=\xi^{n}$. For $n=0$ this holds as $W(0)=\{\langle 1\rangle\}$. Let $n>0$. Recall that $\ell(w)$ denotes the last vertex of a walk $w$. Using Propositions 2.2 and 2.4, the $\xi$-convexity assumption and induction, we get
$h(W(n))$ $=$ ${\textstyle\sum_{w\in W(n)}h(w)}$
$=$ ${\textstyle\sum_{\begin{subarray}{c}w^{\prime}\in W(n-1)\\ \ell(w^{\prime})\in V(h)\end{subarray}}h(w^{\prime})\sum_{v\in\mathbb{N}\setminus\{\ell(w^{\prime})\}}h(\{\ell(w^{\prime}),\,v\})}$
$=$ ${\textstyle\sum_{w^{\prime}\in W(n-1)}h(w^{\prime})\cdot\xi=\xi^{n-1}\cdot\xi=\xi^{n}\,.}$
We can add and remove the condition $\ell(w^{\prime})\in V(h)$ without affecting the sum because walks $w^{\prime}$ not satisfying it have zero weight. $\Box$
3 Extensions of Pólya’s theorem with $v=1$
We begin with the case $v=1$ when the vertex to be visited by walks coincides with their starting vertex. In Theorems 3.23.7 we establish the first group of generalizations of Pólya’s theorem.
For that end, we need some generating functions. Let $n\in\mathbb{N}_{0}$, $v\in\mathbb{N}$, and $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a light weight. Recall that $W(n)$ is the set of walks of length $n$ in $K_{\mathbb{N}}$ starting at $1$, and that $W(v,n)\subset W(n)$ is the subset of walks visiting $v$ in some step $i>0$. We are interested in the generating function
$A_{h}(x)=\sum_{n\geq 0}a_{n}^{h}x^{n}=\sum_{n\geq 0}h(W(1,\,n))x^{n}$}
and define three more generating functions
$B_{h}(x)$ $=$ ${\textstyle\sum_{n\geq 0}b_{n}^{h}x^{n}=\sum_{n\geq 0}h(W_{b}(n))x^{n}\,,}$
$C_{h}(x)$ $=$ ${\textstyle\sum_{n\geq 0}c_{n}^{h}x^{n}=\sum_{n\geq 0}h(W_{c}(n))x^{n}\,\text{ and}}$
$D_{h}(x)$ $=$ ${\textstyle\sum_{n\geq 0}d_{n}^{h}x^{n}=\sum_{n\geq 0}h(W(n))x^{n}\,,}$
all in $\mathbb{C}[[x]]$. Here $W_{b}(n)$ are the walks of length $n$ in $K_{\mathbb{N}}$ that start and end at $1$, and $W_{c}(n)$ are the walks of length $n$ in $K_{\mathbb{N}}$ that start and end at $1$, and avoid $1$ between. The coefficients $a_{n}^{h}$, $\dots$, $d_{n}^{h}$ are correctly defined because they are sums of subseries of the absolutely convergent series
$d_{n}^{h}=\sum_{w\in W(n)}h(w)\,.$}
The constant coefficients are $a_{0}^{h}=c_{0}^{h}=0$ and $b_{0}^{h}=d_{0}^{h}=1$.
@@ 3 Extensions of Pólya’s theorem with $v=1$
Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a light weight and $A_{h}(x)$, $B_{h}(x)$, $C_{h}(x)$ and $D_{h}(x)$ be the above generating functions. The following relations hold.
@@
Proof. 1. We show that $a_{n}^{h}=\sum_{j=0}^{n}c_{j}^{h}\cdot d_{n-j}^{h}$ for every $n\in\mathbb{N}_{0}$. For $n\leq 1$ it holds, then $a_{n}^{h}=c_{n}^{h}=0$. For $n\geq 2$ the set $W(1,n)$ is countable and we split every walk in it at the first revisit of $1$. We get a map
$F\colon W(1,\,n)\to\bigcup_{j=0}^{n}W_{c}(j)\times W(n-j)$}
@@
$w=\langle u_{0},\,u_{1},\,\dots,\,u_{j},\,\dots,\,u_{n}\rangle\in W(1,\,n)\,$,
where $u_{0}=u_{j}=1$, $j>0$ and $u_{i}\neq 1$ for $i\in[j-1]$, we set
$F(w)=\langle w_{1},\,w_{2}\rangle\ \ (\in W_{c}(j)\times W(n-j))\,$,
with $w_{1}=\langle u_{0},u_{1},\dots,u_{j}\rangle$ and $w_{2}=\langle u_{j},u_{j+1},\dots,u_{n}\rangle$. It is easy to see that $F$ is a bijection. $F$ is also weight-preserving (WP): for every weight $h\colon\mathbb{N}_{2}\to\mathbb{C}$ and value $F(w)=\langle w_{1},w_{2}\rangle$, we have $h(w)=h(w_{1})h(w_{2})$. If $X$ is a set of pairs of walks in $K_{\mathbb{N}}$, we write $h(X)$ for the sum of the series
@@
$a_{n}^{h}$ $=$ $h(W(1,\,n))\stackrel{{\scriptstyle\mathrm{WP}}}{{=}}h(F[W(1,\,n)])$
@@ 2. We show that
$B_{h}(x)=\frac{1}{1-C_{h}(x)}=1+\sum_{j\geq 1}(C_{h}(x))^{j}\,.$}
We recall that $b_{0}^{h}=1$ and $c_{0}^{h}=0$ and show that for every $n\in\mathbb{N}$,
$b_{n}^{h}=\sum_{j=1}^{\infty}\sum_{\begin{subarray}{c}n_{1},\,\dots,\,n_{j}\in\mathbb{N}\\ n_{1}+\dots+n_{j}=n\end{subarray}}c_{n_{1}}^{h}c_{n_{2}}^{h}\dots c_{n_{j}}^{h}\,.$}
For odd $n$ it is true: for odd $n$ we have $b^{h}_{n}=c^{h}_{n}=0$, and in every inner summand at least one index $n_{i}$ is odd. For even $n\geq 2$ the set $W_{b}(n)$ is countable. We split every walk in it at the visits of $1$ and get a map
$F\colon W_{b}(n)\to\bigcup_{j=1}^{\infty}\bigcup_{\begin{subarray}{c}n_{1},\,\dots,\,n_{j}\in\mathbb{N}\\ n_{1}+\dots+n_{j}=n\end{subarray}}W_{c}(n_{1})\times\cdots\times W_{c}(n_{j})$}
@@
$w=\langle u_{0},\,\dots,\,u_{m_{1}},\,\dots,\,\dots,\,u_{m_{j-1}},\,\dots,\,u_{n}\rangle\in W_{b}(n)\,$,
where $0=m_{0}<m_{1}<\dots<m_{j}=n$ and $u_{k}=1$ iff $k=0$ or $k=m_{i}$ for some $i\in[j]$, we set
$F(w)=\langle w_{1},\,w_{2},\,\dots,\,w_{j}\rangle\ \ (\in W_{c}(n_{1})\times W_{c}(n_{2})\times\dots\times W_{c}(n_{j}),\,n_{i}=m_{i}-m_{i-1})\,$,
with $w_{i}=\langle u_{m_{i-1}},u_{m_{i-1}+1},\dots,u_{m_{i}}\rangle$, $i\in[j]$. It is easy to see that $F$ is a bijection and that it is weight-preserving, as in part 1. If $X$ is a set of tuples of walks, we write $h(X)$ for the sum of the series
$\sum_{\langle w_{1},\,\dots,\,w_{k})\in X}h(w_{1})\dots h(w_{k})\,$,}
if it absolutely converges.
Let $h$ be a light weight and $n\geq 2$ be even. Using the map $F$, we get
$b_{n}^{h}$ $=$ $h(W_{b}(n))\stackrel{{\scriptstyle\mathrm{WP}}}{{=}}h(F[W_{b}(n)])$
@@
We label our fourteen generalizations of Pólya’s theorem for better orientation by the triples
$\langle x,\,y,\,z\rangle\in\{v=1,\,v\neq 1\}\times\{\mathrm{gen},\,\mathrm{con}\}\times\{U(1)_{1},\,U(1)_{2},\,U(1)_{1}=\infty\}\,$,
where “gen” refers to general edge weights $h\colon\mathbb{N}_{2}\to\mathbb{C}$ and “con” to $\xi$-convex ones. For $U(1)_{1}$ and $U(1)_{2}$ see Definition 2.12, and for $U(1)_{1}=\infty$ see Definition 2.14. The generating functions $A_{h}(x)$, $B_{h}(x)$, $C_{h}(x)$, and $D_{h}(x)$ are defined at the beginning of this section.
Theorem 3.2 (1: $v=1$, gen, $U(1)_{1}$)
If $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is a light weight such that $A_{h}(1)_{1}$, $B_{h}(1)_{1}$, $C_{h}(1)_{1}$ and $D_{h}(1)_{1}$ exist, then $B_{h}(1)_{1}\neq 0$ and
@@
Theorem 3.3 (2: $v=1$, gen, $U(1)_{2}$)
If $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is a light weight such that $\sum_{n\geq 1}|b_{n}^{h}|<1$ and $D_{h}(1)_{2}$ exists, then $B_{h}(1)_{2}\neq 0$, and
@@
$(1-C_{h}(1))_{2}=\frac{1}{B_{h}(1)_{2}},\,\text{ so that }\,C_{h}(1)_{2}=1-\frac{1}{B_{h}(1)_{2}}\,.$}
@@
Theorem 3.4 (3: $v=1$, gen, $U(1)_{1}=\infty$)
If $\xi\in\mathbb{C}_{1}$ and $h\colon\mathbb{N}_{2}\to\xi\mathbb{R}_{\geq 0}$ is a light weight such that $(D_{h})_{1/\xi}(1)_{1}=+\infty$, then
@@
Proof. From $(D_{h})_{1/\xi}(1)_{1}=+\infty$ it follows that $\xi^{-n}d_{n}^{h}>0$ for some $n\in\mathbb{N}$. Hence there is a vertex $u\in\mathbb{N}\setminus\{1\}$ with $\xi^{-1}h(\{1,u\})>0$. The walk $\langle 1,u,1\rangle$ shows by the subseries bound principle that $\xi^{-2}c_{2}^{h}>0$. Since by item 1 of Proposition 3.1 we have $a_{n}^{h}=\sum_{j=0}^{n}c_{j}^{h}d_{n-j}^{h}$, we get that
$\xi^{-2-n}a_{n+2}^{h}\geq\xi^{-2}c_{2}\cdot\xi^{-n}d_{n}^{h}$
@@
We proceed to $\xi$-convex weights.
Theorem 3.5 (4: $v=1$, con, $U(1)_{1}$)
1. If $\xi\in\mathbb{C}_{1}$, $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is a $\xi$-convex light weight and if $(B_{h})_{1/\xi}(1)_{1}$ and $(C_{h})_{1/\xi}(1)_{1}$ exist, then $(B_{h})_{1/\xi}(1)_{1}\neq 0$ and
$\lim_{n\to\infty}\xi^{-n}a_{n}^{h}=1-\frac{1}{(B_{h})_{1/\xi}(1)_{1}}\,.$
2. If $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is a 0-convex light weight, and if $B_{h}(1)_{1}$ and $C_{h}(1)_{1}$ exist, then $B_{h}(1)_{1}\neq 0$ and

### v8

@@ 2.9 Generating functions of walks for $\xi$-convex weights
$\sum_{n\geq 0}h(D_{n})x^{n}=\frac{1}{1-\xi x}\ \ (\in\mathbb{C}[[x]])\,.$}
Thus for $\xi=0$ we have $\sum_{n\geq 0}h(D_{n})x^{n}=1$ ($1x^{0}+0x^{1}+0x^{2}+\dots$).
Proof. Let $\xi\in\mathbb{C}$ and $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a light and $\xi$-convex weight. We prove by induction on $n\in\mathbb{N}_{0}$ that $h(D_{n})=\xi^{n}$. For $n=0$ this holds, $h(D_{0})=1$, because $D_{0}=\{\langle 1\rangle\}$. Let $n>0$. Recall that $\ell(w)$ denotes the last vertex of a walk $w$. Using Propositions 2.2 and 2.4, the assumption of $\xi$-convexity and induction, we get
$h(D_{n})$ $=$ ${\textstyle\sum_{\begin{subarray}{c}w\in D_{n-1}\\ \ell(w)\in V(h)\end{subarray}}h(w)\cdot\sum_{v\in\mathbb{N}\setminus\{\ell(w)\}}h(\{\ell(w),\,v\})}$
$=$ ${\textstyle\sum_{w\in D_{n-1}}h(w)\cdot\xi=h(D_{n-1})\cdot\xi=\xi^{n-1}\cdot\xi=\xi^{n}\,.}$
We can add and remove the condition $\ell(w)\in V(h)$ without affecting the sum because walks $w$ not satisfying it have zero weight. $\Box$
3 General Pólya’s theorems for $v=1$
In this section, we generalize the results in Section 1.2 to weights $h\colon\mathbb{N}_{2}\to\mathbb{C}$. We first introduce three auxiliary generating functions $B_{h}(x)$, $C_{h}(x)$ and $D_{h}(x)$. In Sections 3.2 and 3.3, we obtain general Pólya’s theorems for $v=1$ with general and $\xi$-convex weights, respectively.
3.1 Auxiliary generating functions
Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a light weight (Definition 2.1). As we know from Section 1.4, we are interested in the main generating function
$A_{h}(x)=A_{1,\,h}(x)=\sum_{n\geq 0}h(A_{n})x^{n}\ \ (\in\mathbb{C}[[x]])$}
for the set $A$ of walks in $K_{\mathbb{N}}$ starting at $1$ and revisiting it later; $h(A_{n})$ is the sum of the series
$\sum_{w\in A_{n}}h(w)=\sum_{\begin{subarray}{c}w\in A\\ s(w)=n\end{subarray}}h(w)\,.$}
As in Section 1.2, we consider the set $B$ of closed walks in $K_{\mathbb{N}}$ starting and ending at $1$, the set $C\subset B$ of closed walks with inner vertices different from $1$, and the set $D$ of all walks in $K_{\mathbb{N}}$ starting at $1$. We define three auxiliary generating functions
$B_{h}(x)=\sum_{n\geq 0}h(B_{n})x^{n},\ C_{h}(x)=\sum_{n\geq 0}h(C_{n})x^{n},\ D_{h}(x)=\sum_{n\geq 0}h(D_{n})x^{n}\,.$}
The coefficients $h(A_{n})$, $h(B_{n})$ and $h(C_{n})$, $n\in\mathbb{N}_{0}$, are correctly defined because we assume that the coefficients $h(D_{n})$ are defined ($h$ is a light weight). We have $h(A_{0})=h(C_{0})=0$ and $h(B_{0})=h(D_{0})=1$. Let $F\colon X\to Y$ be a map and $Z$ be any set. In the next proof we use the notation
$F[Z]=\{F(x)\colon\;x\in X\cap Z\}\ \ (\subset Y)\,.$
@@ 3.1 Auxiliary generating functions
Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a light weight and $A_{h}(x)$, $B_{h}(x)$, $C_{h}(x)$ and $D_{h}(x)$ be the above generating functions. The following identities hold in $\mathbb{C}[[x]]$.
@@
Proof. 1. We show that $h(A_{n})=\sum_{j=0}^{n}h(C_{j})\cdot h(D_{n-j})$ for every $n\in\mathbb{N}_{0}$. For $n\leq 1$ it holds because then $h(A_{n})=h(C_{n})=0$. For $n\geq 2$ the set $A_{n}$ is infinite and countable. We split every walk in it at the first revisit of $1$ and get a map
$F\colon A_{n}\to\bigcup_{j=0}^{n}C_{j}\times D_{n-j}$}
@@
$w=\langle u_{0},\,u_{1},\,\dots,\,u_{j},\,\dots,\,u_{n}\rangle\in A_{n}\,$,
where $u_{0}=u_{j}=1$, $j>0$ and $u_{i}\neq 1$ for every $i=1,2,\dots,j-1$, we set
$F(w)=\langle w_{1},\,w_{2}\rangle\ \ (\in C_{j}\times D_{n-j})\,$,
where $w_{1}=\langle u_{0},u_{1},\dots,u_{j}\rangle$ and $w_{2}=\langle u_{j},u_{j+1},\dots,u_{n}\rangle$. It is easy to see that $F$ is a bijection. $F$ is also weight-preserving (WP): for every weight $h\colon\mathbb{N}_{2}\to\mathbb{C}$ and value $F(w)=\langle w_{1},w_{2}\rangle$, we have $h(w)=h(w_{1})h(w_{2})$. If $X$ is a set of pairs of walks in $K_{\mathbb{N}}$, we denote by $h(X)$ the sum of the series
@@
$h(A_{n})$ $\stackrel{{\scriptstyle\mathrm{WP}}}{{=}}$ $h(F[A_{n}])$
@@ 2. We show that
$B_{h}(x)=\frac{1}{1-C_{h}(x)}=1+\sum_{j\geq 1}(C_{h}(x))^{j}\ \ (\in\mathbb{C}[[x]])\,.$}
We recall that $h(B_{0})=1$ and show that for every $n\in\mathbb{N}$,
$h(B_{n})=\sum_{j\geq 1}\sum_{\begin{subarray}{c}n_{1},\,n_{2},\,\dots,\,n_{j}\in\mathbb{N}\\ n_{1}+n_{2}+\dots+n_{j}=n\end{subarray}}h(C_{n_{1}})h(C_{n_{2}})\dots h(C_{n_{j}})\,.$}
For $n=1$ it holds because $h(B_{1})=h(C_{1})=0$. For $n\geq 2$ the set $B_{n}$ is infinite and countable. We split every walk in it at the visits of $1$ and get a map
$F\colon B_{n}\to\bigcup_{j\geq 1}\bigcup_{\begin{subarray}{c}n_{1},\,n_{2},\,\dots,\,n_{j}\in\mathbb{N}\\ n_{1}+n_{2}+\dots+n_{j}=n\end{subarray}}C_{n_{1}}\times C_{n_{2}}\times\cdots\times C_{n_{j}}$}
@@
$w=\langle u_{0},\,\dots,\,u_{m_{1}},\,\dots,\,\dots,\,u_{m_{j-1}},\,\dots,\,u_{n}\rangle\in B_{n}\,$,
where $0=m_{0}<m_{1}<\dots<m_{j}=n$ and $u_{k}=1$ iff $k\in\{0,m_{1},m_{2},\dots,m_{j}\}$, we set
$F(w)=\langle w_{1},\,w_{2},\,\dots,\,w_{j}\rangle\ \ (\in C_{n_{1}}\times C_{n_{2}}\times\dots\times C_{n_{j}},\,n_{i}=m_{i}-m_{i-1})\,$,
where $w_{i}=\langle u_{m_{i-1}},u_{m_{i-1}+1},\dots,u_{m_{i}}\rangle$, $i=1,2,\dots,j$. It is easy to see that $F$ is a bijection and that it is weight-preserving (WP), as in part 1. If $X$ is a set of $k$-tuples of walks, we denote by $h(X)$ the sum of the series
$\sum_{\langle w_{1},\,w_{2},\,\dots,\,w_{k})\in X}h(w_{1})h(w_{2})\dots h(w_{k})\,$,}
if the series absolutely converges.
Let $h$ be a light weight and $n\geq 2$. Using the map $F$, we get
$h(B_{n})$ $\stackrel{{\scriptstyle\mathrm{WP}}}{{=}}$ $h(F[B_{n}])$
@@
We already mentioned in Section 1.5 that the previous proof differs from all related proofs in the first phase of the symbolic method because we work with infinite sets of combinatorial objects,
3.2 Pólya’s theorems for $v=1$ and general weights
We obtain general Pólya’s theorems of the form indicated in the subsection heading. For the labels $\exists\,U(1)_{1}$ and $\exists\,U(1)_{2}$ of the theorems see Definition 2.12, and for the label $U(1)_{1}=\infty$ see Definition 2.14. The generating functions $A_{h}(x)$, $B_{h}(x)$, $C_{h}(x)$, and $D_{h}(x)$ are defined in the previous subsection.
Theorem 3.2 ($\exists\,U(1)_{1}$)
Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a light weight. If $A_{h}(1)_{1}$, $B_{h}(1)_{1}$, $C_{h}(1)_{1}$ and $D_{h}(1)_{1}$ exist, then $B_{h}(1)_{1}\neq 0$ and
@@ 3.2 Pólya’s theorems for $v=1$ and general weights
Theorem 3.3 ($\exists\,U(1)_{2}$)
Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a light weight. If $\sum_{n\geq 1}|h(B_{n})|<1$ and $D_{h}(1)_{2}$ exists, then $B_{h}(1)_{2}\neq 0$ and
@@
$(1-C_{h})(1)_{2}=\frac{1}{B_{h}(1)_{2}},\,\text{ so that }\,C_{h}(1)_{2}=1-\frac{1}{B_{h}(1)_{2}}\,.$}
@@
Theorem 3.4 ($U(1)_{1}=\infty$)
Let $\xi\in\mathbb{C}_{1}$ and $h\colon\mathbb{N}_{2}\to\xi\mathbb{R}_{\geq 0}$ be a light weight. If $(D_{h})_{1/\xi}(1)_{1}=+\infty$, then
@@
Proof. From $(D_{h})_{1/\xi}(1)_{1}=+\infty$ it follows that $\xi^{-n}h(D_{n})>0$ for some $n$ in $\mathbb{N}$. Hence there is a vertex $u\in\mathbb{N}\setminus\{1\}$ such that $\xi^{-1}h(\{1,u\})>0$. The walk $\langle 1,u,1\rangle$ shows by the subseries bound that $\xi^{-2}h(C_{2})>0$. Since by item 1 of Proposition 3.1 we have $h(A_{n})=\sum_{j=0}^{n}h(C_{j})h(D_{n-j})$, we get that
$\xi^{-2-n}h(A_{n+2})\geq\xi^{-2}h(C_{2})\cdot\xi^{-n}h(D_{n})$
@@
3.3 Pólya’s theorems for $v=1$ and $\xi$-convex weights
We proceed to $\xi$-convex weights and begin with twisted nonnegative ones. The next theorem generalizes Section 1.2.
Theorem 3.5 ($\exists\,U(1)_{1}$ and $h\geq 0$)
Let $\xi\in\mathbb{C}_{1}$ and $h\colon\mathbb{N}_{2}\to\xi\mathbb{R}_{\geq 0}$ be a $\xi$-convex light weight. Then, with $(B_{h})_{1/\xi}(1)_{1}\in[1,+\infty)\cup\{+\infty\}$, we have
$\lim_{n\to\infty}\xi^{-n}h(A_{n})=1-\frac{1}{(B_{h})_{1/\xi}(1)_{1}}\,.$
For $(B_{h})_{1/\xi}(1)_{1}=+\infty$ the limit is $1$, and else it lies in $(0,1)$.
Proof. Proposition 2.31 gives $D_{h}(x)=\frac{1}{1-\xi x}$. Since $B_{n},C_{n}\subset D_{n}$ for every $n\in\mathbb{N}_{0}$, by the subseries bound we have $\xi^{-n}h(B_{n}),\xi^{-n}h(C_{n})\in[0,1]$. Now we argue as in Section 1.2. The last claim that $(B_{h})_{1/\xi}(1)_{1}>1$ follows from the assumption of $\xi$-convexity: there is a $u\in\mathbb{N}\setminus\{1\}$ such that $\xi^{-1}h(\{1,u\})>0$, and hence $\xi^{-2}h(B_{2})>0$. $\Box$
Theorem 3.6 ($\exists\,U(1)_{1}$)
1. Let $\xi\in\mathbb{C}_{1}$ and $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a $\xi$-convex light weight. If $(B_{h})_{1/\xi}(1)_{1}$ and $(C_{h})_{1/\xi}(1)_{1}$ exist, then $(B_{h})_{1/\xi}(1)_{1}\neq 0$ and
$\lim_{n\to\infty}\xi^{-n}h(A_{n})=1-\frac{1}{(B_{h})_{1/\xi}(1)_{1}}\,.$
2. Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a 0-convex light weight. If $B_{h}(1)_{1}$ and $C_{h}(1)_{1}$ exist, then $B_{h}(1)_{1}\neq 0$ and